# Residual Diffusion Modeling for Km-scale Atmospheric Downscaling

### v3

@@ 4 Examining sample diversity of CorrDiff
In order to examine the realism of individual samples from CorrDiff and their quality compared with the target data figure shows an animation of the target data (left) Reg (middle) and 20 generated samples of the CorrDiff prediction of maximum radar reflectivity.
@@
[width=autoplay,loop]1000000034
Figure S3: Comparison of many samples of maximum radar reflectivity (right) with the prediction from the regression (Reg) and the truth (WRF) for several cloud regimes.
5 Comparison of simulated Typhoons with historical records
@@
Although our out-of-sample data is limited to 2021, we can compare CorrDiff simulated typhoons in the CWA region with historical records. The Japan Meteorological Agency best track data (JMA tracks) [4] includes the maximum windspeed (intensity) and radius of maximum windspeed (size) of typhoons in west pacific for several decades. During the years 1980 to 2020 for (which we have ERA5 data available) and within the CWA domain, we identified 648 instances of typhoons with intensities of 30 m/s or greater in this time period. Panels (a) and (b) of figure S4 display the storm size and intensity, respectively, revealing the expected correction for ERA5 typhoons achieved through the application of CorrDiff. One imperfection of CorrDiff is that it shrinks all storms, including those with the correct size or those that were too small in the ERA5 input data (panel a). The main benefit is an improvement in the windspeeds, removing most of the error between the ERA5 and the observed records for windspeeds up to 50 m/s (panel b); stronger storms have room for improvement. CorrDiff generates a five-fold increase in the probability of hurricane-force winds exceeding 33 m/s (panel c). To the extent that JMA tracks can serve as ground truth, such distribution shift has significant societal implications as these low-probability, high-impact events represent a substantial portion of the overall risk, underscoring the importance of accurate modeling and forecasting.
Figure S4: Comparison between ERA5, CorrDiff (from ERA5) and observed records for all typhoons found in the domain with windspeed of 30 m/s or more in the years 1980-2020. Panel (a): abscissa: mean (dot) and one standard deviation (bars) of predicted radius of maximum winds for the ERA5 (red) and CorrDiff (orange) storms, grouped by their observed values; ordinate: corresponding observed values. Panel (b) same as (a) but the maximum windspeed. Panel (c): The PDFs of all instances with typhoon of windspeed 30 m/s or greater in the years 1980-2020 in the CWA domain.
@@
6 Energy efficiency and latency of CorrDiff donwscaling inference compared to WRF simulations
@@
Comparing the performance of a statistical downscaling like CorrDiff with a dynamical downscaling like the WRF-CWA is not straightforward. The former produces maps of a subset of channels at time $t+1$ from a low resolution global forecast at $t+1$, while the latter produces maps of the full state vector at time $t+1$ from a low resolution global forecast at $t$ using a numerical time-stepper where each time step is of the order of seconds.
@@
Nonetheless, given that a global model like GFS is available, we can compare the compute time and energy required to obtain high resolution 1 hourly maps in Taiwan from the two approaches. Doing so, we compare the speed of CorrDiff against the operational WRF run by CWA on their respective hardware. The CWA-WRF is run on Fujitsu FX-100 system, where each node is equipped with 32 SPARC64 Xifx CPU cores. A 13 hour deterministic CWA-WRF forecast (excluding data assimilation) is run on 928 CPU cores (across 29 nodes with a maximum system memory of 6.9GB per node) and takes about 20 minutes. CorrDiff inference is run on a single NVIDIA RTX 6000 Ada Generation GPU, which takes 4.1 sec per downscaling sample. Given a global model 1 hour leadtime forecast, the CorrDiff statistical downscaling on a single GPU is about $22$ times faster than the dynamic downscaling that runs CWA-WRF on 928 CPUs. Moreover, CorrDiff it is about $1130$ times more energy efficient. We note that our current implementation of CorrDiff is far from optimized and does not utilize GPU parallelization and batching for generating many samples independently. A simple change to mixed-precision (float 16) gains almost another factor 2 speedup, without any degradation of the results, leading to about 38×38$\times$38 × speed up and about 1,800×1,800$\times$1 , 800 × energy efficiency compared with WRF-CWA. Furthermore, given global model data (which both systems need) CorrDiff can be run for the 13h on 13 GPUs obtaining about a 13x speedup for the 13h forecast used, over the above results (but with the same energy efficiency). We believe that GPU parallelization and batched inference will likely lead to even more significant speed-up and energy improvements.
@@
CorrDiff, float32 1 GPU 4.1 300 1.23
CorrDiff, float16 1 GPU 2.4 300 0.72
Table S3: A comparison of running the WRF model on the CWA system with CorrDiff inference on a single NVIDIA RTX 6000 Ada Generation GPU. Latency is given per Forecast Hour (FH) and Power is given in Joule/sec (W) per a single hardware unit (a CPU or a GPU), while Energy is for the entire forecast system (928 CPU for WRF-CWA) per FH.

### v4

@@ 4 Examining sample diversity of CorrDiff
In order to showcase the realism of individual samples from CorrDiff and their quality compared with the target data, Fig. S4 shows an animation of the target data (left), UNet (middle), and 20 generated samples of the CorrDiff prediction (right) of maximum radar reflectivity.
@@
[width=autoplay,loop]10animation/000000159
Figure S4: Comparative analysis of maximum radar reflectivity across multiple samples: Diffusion model predictions versus UNet regression and WRF target simulations for diverse cloud regimes. This figure is presented as an animated visualization, viewable as a video in Adobe-compatible formats. The full animation can be accessed and downloaded from https://figshare.com/ndownloader/files/48060031.
5 Pooled metrics
Due to the stochastic nature of km atmospheric fields, and specifically radar reflectivity, we complement the main text with a pooled counterpart of the metrics reported in 1. In S3 the MAE and CRPS are computed from data data pooled over 14 points, which is roughly the resolution of the ERA5 data used for conditioning.
Radar t2m u10m v10m
CorrDiff (CPRS) 1.64 0.5 0.73 0.82
CorrDiff (MAE) 2.0 0.59 0.88 0.98
UNet (MAE) 1.98 0.53 0.89 0.99
RF (MAE) 2.91 0.62 0.91 1.01
ERA5 (MAE) - 0.83 0.97 1.05
Table S3: pooled CRPS counterpart table for 1, where the prediction and target from each model are pooled on 28 grid boxes.
Here as well, CorrDiff CRPS is superior to the MAE for all baselines, followed by the UNet, RF and ERA5 interpolation. This implies that the results presented in the main text are a valid reflection of the model’s performance compared with the baselines.
6 Additional Case Study Analysis
To complement the example we have in the main text, we provide here an additional analysis of the case studies. For weather fronts, we analyze the collocation of the front with the reflectivity (clouds and rain), and for typhoons, we examine two different typhoons as well as historical typhoons for which there is no target data.
6.1 Additional Weather Front Analysis
Building on the analysis of the front in the main text (see Fig. 5), we further examine the coherence of synthesized reflectivity in two frontal events: one from 2022 (mentioned above) and another from 2023. Figures S5 and S6 compare the reflectivity maps for the target and CorrDiff along with the cross-section of the along-front winds.
This analysis shows that although CorrDiff does not always sharpen the fronts to meet the target data, it syntheses the radar reflectivity consistently with the other variables and respects the frontal boundary by keeping the warm sector (of negative along front winds) cloud-free. Here we choose to indicate the frontal boundary with the along front wind. In all frontal events we examined, except the 2022 event in the main paper, fronts are moving fast and have large scale components of across front temperature gradient, and winds, which complicates tracking them in these channels.
Figure S5: Analyzing the reflectivity synthesized during a cold front. Left to right: radar reflectivity maps of CorrDiff and the target data, followed by the transects of reflectivity and of the along front wind. along the dashed line in the adjacent maps. Top to bottom:2022-02-13 17:000:00, 2022-02-13 19:00:00, 2022-02-13 21:00:00 and 2022-02-13 23:00:00 UTC.
Figure S6: Same as S5 but for time: 2023-02-13 08:00:00, 2023-02-13 10:00:00, 2023-02-13 12:00:00 and 2023-02-13 14:00:00 UTC.
6.2 Additional Typhoon Analysis
As additional analysis we first analyze two out-of-sample typhoons and then analyze a large number of typhoons without CWA target data by comparing their diagnosed properties to observed records.
Figures S7 and S8 below show analyses of typhoons Chanthu (2021) and Haikui (2023) respectively. The former is in our out-of-sample year and the latter was received in an additional data from CWA specifically for this analysis. The performance of ERA5, and consequently CorrDiff, differs substantially between these two typhoons. Chanthu (2021) presents an exceptional challenge for low resolution models like the one underlying ERA5, due to its meridional trajectory along Taiwan’s coast at a distance of about 100-200 km. In such a trajectory the impact of Taiwan ’s steep topography (up to 3km elevation) depends on the size of the simulated typhoon. The small Chanthu (2021) simulated by the WRF model, with a radius of maximum winds of about 25km, is effectively over ocean. Conversely, the larger Chanthu (2021) in ERA5, with radius of maximum winds of about 100km, is significantly affected by the topography.
As a result, the ERA5 Chanthu (2021) can have less than a third of the intensity its WRF counterpart, see the first and third columns in S7. It often does not have a closed windspeed contour above $10ms^{-1}$, while the WRF Chanthu (2021) displays a coherent structure at $50ms^{-1}$ contours. Such a gap between the target (WRF) and the condition (ERA5) is hard for CorrDiff to overcome, and indeed the model fails to recover the intensity of the typhoon for most days. Only on 2023-09-03 12:00:00 UTC (top row of S7), when the typhoon has passed the island, does CorrDiff provide an improvement over ERA5.
Figure S7: Windspeed maps and axisymmetric windspeed of typhoon Chanthu (2021). Left to right: windspeed in ERA5, CorrDiff, target (WRF model) and their axisymmetric profiles. The red ‘+‘, orange diamond and black dot show the storm center for the ERA5, CorrDiff, WRF respectively. Time is increasing from bottom to top: 2021-09-02 12:00:00 UTC, 2021-09-02 18:00:00 UTC, 2021-09-03 00:00:00 UTC, 2021-09-03 06:00:00 UTC, 2021-09-03 12:00:00.
Unlike Chanthu (2021), typhoon Haikui (2023) has a zonal trajectory leading to landfall in southern Taiwan. CorrDiff improves upon ERA5 by correcting about 50% of the intensity error and contracts the typhoon’s radius of maximum winds.
Figure S8: Same as S7 but for typhoon Haikui (2023). Time is increasing from bottom to top: 2023-09-02 18:00:00 UTC, 2023-09-02 21:00:00 UTC, 2023-09-03 00:00:00 UTC, 2023-09-03 03:00:00 UTC, 2023-09-03 06:00:00 UTC.
@@ 6.2 Additional Typhoon Analysis
Further analysis compares CorrDiff-simulated typhoons in the CWA region with historical records to evaluate typhoons for which no target data exists. The Japan Meteorological Agency best track data (JMA tracks) [4] includes the maximum windspeed (intensity) and radius of maximum windspeed (size) of typhoons in the West Pacific for several decades. We identified 648 instances of typhoons with intensities of $30ms^{-1}$ or greater within the CWA domain from 1980 to 2020. Panels (a) and (b) of Fig. S9 display the storm size and intensity, respectively, revealing the expected correction for ERA5 typhoons achieved through the application of CorrDiff downscaling. One limitation of CorrDiff is that it reduces the size of all storms, including those with the correct size or those already too small in the ERA5 input data (panel a). The main benefit is improved windspeeds, removing most of the error between the ERA5 and the observed records for windspeeds up to $50ms^{-1}$ (panel b); though stronger storms have room for improvement. By correcting ERA5 toward JMA tracks, CorrDiff generates a five-fold increase in the probability of windspeed values exceeding hurricane-force winds (i.e., $33ms^{-1}$), see panel c. To the extent that JMA tracks can serve as ground truth, such distribution shift has significant societal implications, as these low-probability, high-impact events represent a substantial portion of the overall risk.
The UNet alone presents a less attractive alternative for downscaling typhoons, as its maximum axisymmetric windspeed is consistently positioned between the ERA5 and the CorrDiff values. CorrDiff offers a meaningful improvement over the UNet in both the maximum of the axisymmetric wind speed (panel b) and the wind speed maxima (panel c).
Figure S9: Comparison of predicted vs. observed (a) radius (km) and (b) value ($ms^{-1}$) of maximum axisymmetric windspeed ($ms^{-1}$) for all typhoons found in the domain with windspeed of $30ms^{-1}$ or more spanning the years 1980-2020. Predictions are compared via the mean (dot) and one standard deviation (bars) of all predictions falling within each observational bin of the reference JMA observations. Panel (c) shows the log(PDF) of the predicted windspeed flattened across all of these typhoon instances.
@@
6.3 Energy efficiency and latency of CorrDiff downscaling inference compared to WRF simulations
@@
Comparing the performance of statistical downscaling like CorrDiff with dynamical downscaling like the WRF-CWA is challenging due to their different approaches and outcomes. Statistical downscaling produces a high-resolution state for a subset of channels at time $t$ from a different (potentially larger) set of channels at lower resolution at the same time. Dynamical downscaling produces a full, high-resolution, state vector at time $t+\Delta t$ from both high and low resolution state vector at $t$ using a numerical time-stepper. The time step ($\Delta t$) here is constrained by numerical stability and can be of the order of seconds. Thus, to produce the prediction at +1h, the dynamical downscaling model might run hundreds of autoregressive steps while the statistical downscaling will make a single inference.
Nonetheless, from a utility perspective, both approaches are used for producing a high resolution state at a given time. Thus, we can compare the latency and energy required to obtain a single high resolution prediction in Taiwan from the two approaches. We compare the speed of CorrDiff against the operational WRF run by CWA on their respective hardware.
@@
The CWA-WRF is run on Fujitsu FX-100 system, with each node equipped with 32 SPARC64 Xifx CPU cores. A 13-hour deterministic CWA-WRF forecast (excluding data assimilation) is run on 928 CPU cores (across 29 nodes with a maximum system memory of 6.9GB per node) and takes about 20 minutes. CorrDiff inference is run on a single NVIDIA H100 Generation GPU, which takes 0.18 sec per downscaling sample. Given a global model 1-hour lead time forecast, the CorrDiff statistical downscaling on a single GPU is about $500$ times faster than the dynamic downscaling that runs CWA-WRF on 928 CPUs. Moreover, CorrDiff is about $10,000$ times more energy efficient. Since the individual samples are computed independently, conditioned on given global model data (which both systems need), CorrDiff can be run for the 13 hours on 13 GPUs, thus obtaining about a 13x speedup for the 13-hour forecast over the above results (but with the same energy efficiency). These results ignore the additional compute needed for the regional DA in CWA-WRF, which is absent in CorrDiff (and likely impacts its performance). The regional DA in CWA, depending on the method used, can increase the compute for the CWA-WRF by a factor of 1.5 or 2 [15].
@@
CorrDiff 1 GPU 0.18 700 0.126
Table S4: A comparison of running the WRF model on the CWA system with CorrDiff inference on a single NVIDIA H100 GPU. Latency is given per Forecast Hour (FH) and Power is given in Joule/sec (W) per a single hardware unit (a CPU or a GPU), while Energy is for the entire forecast system (928 CPU for WRF-CWA) per FH.
6.4 Statistical significance of the CRPS metrics
Figure S10: Boostrapping analysis of skill improvements presented in Table 1. The error bar shows a 95% confidence interval for the mean obtained by bootstrapping.
We further diagnose the statistical significance of the metrics presented in Table 1. We show this significance both graphically and using hypothesis testing.
By evaluating our models on the same set of times, we can take advantage of paired statistical tests, which provide much more power. Given the scores for two models $x_{i}$ and $y_{i}$ evaluated over times $i$, we find that $Var(x_{i}-y_{i})\ll Var(x_{i})$. So even though the scores may vary in time, the improvements are robust. Figure S10 shows the improvement in CRPS relative to the RF baseline. In all cases, the improvements of our models (CorrDiff and UNet) are larger than the error bars.
We further elaborate on the significance with formal hypothesis testing for the radar field. The standard non-parametric test for assessing whether $x_{i}>y_{i}$ is the Wilcoxon signed-rank test for $d_{i}=(x_{i}-y_{i})$. An alternative test is the binomial test of $x_{i}>y_{i}$. In both cases, the p-values comparing the RF baseline with CorrDiff are smaller than $10^{-30}$. This remarkably low p-value may seem surprising, but CorrDiff has lower CRPS in 205 out of 205 times. The likelihood this occurring by random chance is vanishingly small.